OpenStudy (anonymous):

integration of sin2x.cos3x dx

OpenStudy (anonymous):

use the trig id: \[\sin(A)\cos(B) = \frac{ 1 }{ 2 }[\sin(A+B) + \sin(A-B)]\] A = 2x B = 3x

OpenStudy (anonymous):

thanks got the answer